OpenStudy (anonymous):

Solve. If the equation has no solution, Write "No Solution" 2[3(x+4)-2(x+1)]=5x+3(1-x) Show work

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

any attempts?

OpenStudy (anonymous):

No.. @MARC_

OpenStudy (anonymous):

alrite

OpenStudy (anonymous):

\[2(3x+12-2x-2)=5x+3-3x\]\[2(3x-2x+12-2)=5x-3x+3\]\[2(x+10)=2x+3\]\[2x+20=2x+3\]\[2x-2x=-20+3\]\[0=-17\]Therefore,no solution. @sabby75488

OpenStudy (anonymous):

Okay Can I send you more please ?! @MARC_

OpenStudy (anonymous):

sure. post in a new thread.